OpenStudy (anonymous):

Relations: in my book, it says that a relation is a subset of the set of ordered pairs of real numbers, XY. What do they mean by that?

OpenStudy (anonymous):

Well, simply put, in set theory a relation is any set which contains some ordered pairs.

OpenStudy (anonymous):

It's much easier to think of it graphically. For the universal set R2, a subset of ordered pairs G is a relation. The relation xGy is a set which contains all coordinate pairs (x, y) which satisfy G (ex. x^2 + y^2 = 1 - all the points along a circle).

OpenStudy (experimentx):

do you know about Cartesian product? http://en.wikipedia.org/wiki/Cartesian_product

OpenStudy (anonymous):

What does that have to do with this?

OpenStudy (anonymous):

@jim_thompson5910 yes, I know what they mean by a set of ordered pairs of real numbers eg. \[\left \{ (1,2) (3,4)(-6,8) \right \}\]

OpenStudy (anonymous):

with spaces in between of course

OpenStudy (experimentx):

hmm ... you missed comma between two elements

OpenStudy (anonymous):

@bronzegoddess did u get what I said earlier?

OpenStudy (anonymous):

thats why said with spaces in between

OpenStudy (anonymous):

@vf321 i didn't understand your graphical explanation

OpenStudy (experimentx):